OpenStudy (lovelycharm):

Solve 6/x-3=3/x for x and determine if the solution is extraneous or not.

OpenStudy (mathmale):

Are you dividing that 6 by x or are you dividing it by (x-3)? Please use parentheses around the divisor(s) as necessary, so that there is no doubt regarding what comprises your divisor.

OpenStudy (lovelycharm):

is it non-extraneous or extraneous i got -3

OpenStudy (lovelycharm):

@mathmale

OpenStudy (tkhunny):

You first must answer the question. 6/x-3=3/x If you mean \(\dfrac{6}{x} - 3 = \dfrac{3}{x}\), then you have written it correctly. If you mean \(\dfrac{6}{x-3} = \dfrac{3}{x}\), then parentheses are NOT optional. You shoudl have written, 6/(x-3)=3/x

OpenStudy (mathmale):

@tkhunny: Many thanks for reinforcing this!